OpenStudy (anonymous):

I could definitely use some help with the Trapezoid, Midpoint, and Simpson's rule (where n=6) for: \[\large9\int\limits_{1}^{4}\sqrt{\ln \left| x \right|}\cdot dx\] I think I must be confused by the instructions for these or something because it's not taking my evaluated decimal answers, six significant figures required. (+1 medal as always for steps shown, ty in advance to taking the time to help)

OpenStudy (anonymous):

\[\triangle x = \frac{b-a}{n} = \frac{4-1}{6}=\frac{1}{2}\]

OpenStudy (anonymous):

And so...

OpenStudy (experimentx):

\[ 9 \times\sum_{n = 1}^{6} \frac 12\left ( \sqrt{ \ln \left (1 + (n-1) \frac 1 2 \right ) } + \sqrt{ \ln \left (1 + (n) \frac 1 2 \right ) } \right ) \frac 12 \]

OpenStudy (anonymous):

Err what? I know the definition of an integral is the sum of areas as the width of each sample goes to zero, but I've never seen what you've written before like that @experimentX

OpenStudy (experimentx):

isn't it asking for numerical integration??

OpenStudy (anonymous):

\[T_n= \frac{\triangle x}{2} (f(x_1)+2f(x_2)+...2f(x_{n-1})+f(x_n)\] \[M_n= \triangle x (f(x_1)+f(x_2)+...+f(x_n)\] \[S_n= \frac{\triangle x}{3} (f(x_1)+4f(x_2)+2f(x_3)+2f(x_4)+...2f(x_{n-2})+4f(x_{n-1})+f(x_n)\]

OpenStudy (anonymous):

True it is, but it's asking not for exact, it's asking for approximated by these things.

OpenStudy (anonymous):

To six decimal places. I keep trying to see if I have an error in my formulas or a typo on the calculator or something

OpenStudy (blockcolder):

For Simpson's Rule, the coefficients go 1-4-2-4-2-...-2-4-1.

OpenStudy (blockcolder):

Also, I'm assuming it's a typo, but for Midpoint Rule, the sample points \(x_i^\star\) should be \(x_i^\star=\frac{1}{2}(x_{i-1}+x_i)\)

OpenStudy (anonymous):

@blockcolder , Yay finally somebody who knows this stuff! :D So for the midpoint rule, each f(x) should be f(\(x_i^*\)) where that sample point what you wrote above? I'll have to try that again. And yeah, clearly Simpson's rule in my notes was wrong then if that's how it's supposed to go. In the example problem in the text they only worked out one with n=5, not any larger.

OpenStudy (anonymous):

\[S_n= \frac{\triangle x}{3} (f(1)+4f(1.5)+2f(2)+4f(2.5)+2f(3)+4f(3.5)+f(4))\] Like this?

OpenStudy (blockcolder):

You can't use Simpson's Rule for n=5, because it's odd. For n=6, Simpson's Rule should be: \[S_n=\frac{\Delta x}{3}(f(x_0)+2f(x_1)+4f(x_2)+2f(x_3)+4f(x_4)+2f(x_5)+f(x_6))\] You're forgetting to include \(x_0\), btw. \(x_0=a\)

OpenStudy (anonymous):

Ok I've gotten everything correct except for the midpoint gentlement

OpenStudy (anonymous):

T\(_n \approx\) 23.32200512 S\(_n \approx\) 23.68778685 Ty for your help so far @blockcolder & @experimentX ;-)

OpenStudy (anonymous):

I guess I need a bit of a better understanding as to what the terms of the midpoint rule should be?

OpenStudy (experimentx):

haven't done midpoint rule yet ... let's see what's on google.

OpenStudy (anonymous):

Meh Google didn't help much, it still giving me errors for some reason. There's a second problem tied to this one which also requires using the midpoint, right-endpoint, and left-endpoint based on a drawing posted, not an actual function.

OpenStudy (anonymous):

Here take a look @experimentX This is the other one the follows this:

OpenStudy (experimentx):

the midpoint rule just says \[ \sum \Delta x f(x + \Delta x/2)\]

OpenStudy (anonymous):

Kind of a pain if you ask me, I tried using half-unit samples and as you can see it says NO!

OpenStudy (anonymous):

Granted that's a separate problem there, but it's the same midpoint rule.

OpenStudy (anonymous):

And you can't just average the right and left endpoints apparently

OpenStudy (anonymous):

Which is what I THOUGHT the "midpoint" was all about, no?

OpenStudy (experimentx):

seems all right ... L2<M2<R2

OpenStudy (experimentx):

the area of that curve must be around 46's

OpenStudy (anonymous):

But as you can see I'm getting a big fat NO from anything in the system regarding Midpoint rule, which I feel logical must mean I'm doing something wrong in my process.

OpenStudy (anonymous):

logically*

OpenStudy (experimentx):

can you give me the function?

OpenStudy (anonymous):

You mean for that image one? No that's literally it. The function is what's visually specified, unfortunately

OpenStudy (experimentx):

seems like a parabola ... what's your answer?? try to reduce the step size to get more accurate result.

OpenStudy (anonymous):

For the one proposed initially for this question/topic where I needed to use 3 different methods, that's as written sqrt(ln|x|)). Reduce to 0.25 samples instead of 0.5?

OpenStudy (anonymous):

For \(\triangle x\) I mean

OpenStudy (anonymous):

I'm currently trying to compare this topic with this one: http://www.physicsforums.com/showthread.php?t=157496

OpenStudy (experimentx):

well ... that gives more accurate result for every numerical integration. For the original problem, Sum[ 9 ( Sqrt[Log[1 + n * 0.5]]), { n, 1 , 5 }] on mathematica gives 41.3457

OpenStudy (experimentx):

Sorry, Sum[ 9 ( Sqrt[Log[1 + n * 0.5]])*0.5, { n, 1 , 5 }] = 20.6728

OpenStudy (anonymous):

Alas it says no, to this: http://www.wolframalpha.com/input/?i=Sum [+9+%28+Sqrt[Log[1+%2B+n+*+0.5]]%29*0.5%2C+{+n%2C+1+%2C+5+}]

OpenStudy (experimentx):

Sorry ... i am not getting accurate results. wait for a while.

OpenStudy (anonymous):

err that link broke

OpenStudy (anonymous):

I can see why the coefficient out front is 9/2 easily enough, let me the terms...

OpenStudy (anonymous):

9*\(\triangle x\)(\(\sqrt{\ln|1.25|}\)+\(\sqrt{\ln|1.75|}\)+\(\sqrt{\ln|2.25|}\)+\(\sqrt{\ln|2.75|}\)+\(\sqrt{\ln|3.25|}\)+\(\sqrt{\ln|3.75|}\))

OpenStudy (experimentx):

this is for right sum Sum[ 9 ( Sqrt[Log[1 + n * 0.5 ]])*0.5, { n, 1 , 6 }] = 25.9712 For mid sum Sum[ 9 ( Sqrt[Log[1 + n * 0.5 + 0.25]])*0.5, { n, 0 , 5}] = 24.1294 For left sum Sum[ 9 ( Sqrt[Log[1 + n * 0.5 ]])*0.5, { n, 0 , 5 }] = 20.6728

OpenStudy (experimentx):

well ... what value did you get for that??

OpenStudy (experimentx):

i think it's better to call it upper sum and lower sum rather than left sum or right sum.

OpenStudy (anonymous):

AH HA! 24.12940953421 It wasn't taking 24.1294

OpenStudy (anonymous):

Freaking computer, like 9 millionths matters...

OpenStudy (experimentx):

online test??

OpenStudy (anonymous):

1. How do I manipulate your syntax-based formula for other functions? 2. Can I give you a medal and you give the other guy a medal? I would love to give credit to both of you :D 3. Not quite, online homework stuff.

OpenStudy (anonymous):

One of the few parts I'm stuck on

OpenStudy (anonymous):

It's only a few points, but I'd rather get as much as I can and understand why, takes pressure off the test

OpenStudy (experimentx):

sure ..where are you stuck??

OpenStudy (anonymous):

Well it's this section, and part of it it I think is because of the computer being a punk and not taking it if it's not EXACTLY what it wanted

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I'll create two new topics @experimentX , so you can get extra medals for your extra effort. It's the same stuff so we're on the same page already, and nobody else seems to want to help out with this mess of calculus so you're GREATLY appreciated. One of the few you and @blockcolder are, and my deepest respects to you both :-)

OpenStudy (experimentx):

well ... just check in for value of x Sum[ function goes first, {value of increment, initial value , final value}] in function place modify it by f(x + n*increment step)* increment step.

OpenStudy (experimentx):

no probs ... it's all right. BTW i didn't know midpoint formula until now. well, i have unusual learning methods.